# PreferThinker: Reasoning-based Personalized Image Preference Assessment

### v1

@@ 1 Introduction
Image preference assessment is essential for both evaluating and aligning generative models with human preferences, playing a key role in applications like content recommendation. However, most existing works (Kirstain et al., 2023; Xu et al., 2023) focus on assessing general preferences, such as text-image alignment and aesthetics, relying on large-scale general preference data. The practical but challenging task of personalized preference assessment remains largely underexplored, where only a limited amount of personalized data is available for each user as prior information to assist in the assessment. As shown in Fig. 1 (a), personalized preference assessment is particularly challenging for two main reasons. First, unlike easily scalable general preference data, where most users share similar assessment criteria, each user’s personalized data is typically limited and distinct from others, making large-scale, user-specific training difficult. Second, personalized preferences are often complex and diverse, covering multiple dimensions such as art style, color, and art medium, which are more challenging than clear general human preferences (e.g. text-image alignment and aesthetics).
Current research on image preference assessment mainly falls into two categories: CLIP-based methods and Multimodal Large Language Model (MLLM)-based methods. CLIP-based methods (Wu et al., 2023b; a) typically fine-tune the CLIP model using large-scale text-image pairs with similarity to assess general preferences. However, they struggle with personalized preferences since the amount of personalized data is typically limited for large-scale training. Besides, simple similarity fails to capture complex individual preferences and lacks interpretability, as it outputs only a numerical score.
MLLM-based methods (Wang et al., 2025c; Xu et al., 2024) achieve interpretable preference assessment by fine-tuning the MLLM on large-scale visual question-answering (VQA) pairs. However, the scarcity of personalized images makes it difficult to obtain sufficient VQA pairs for large-scale training, which limits their effectiveness for personalized preference assessment. Overall, most existing methods are designed for general preferences and therefore struggle to handle personalized preference. Recently, an MLLM-based method named ViPer (Salehi et al., 2024) was proposed for personalized preference assessments. It directly feeds both personalized reference images and candidate images into an MLLM and fine-tunes the model using a supervised score regression strategy. Nevertheless, it only implicitly utilizes reference images to assist in evaluating the candidate images, failing to fully leverage the critical prior information they contain. Moreover, it lacks interpretable reasoning steps that explain how to produce its final score based on the individual reference images.
@@ 2 Related Work
Image Preference Assessment is crucial for both evaluating and aligning generative models with human preferences. While metrics such as FID (Heusel et al., 2017) and CLIP scores (Radford et al., 2021) assess image quality and text-image consistency, they are limited in capturing human preferences. To address this, most works (Kirstain et al., 2023; Liang et al., 2024; Li et al., 2024; Zhang et al., 2024; Ba et al., 2025) employ large-scale preference data to fine-tune CLIP models for assessing general preferences. However, they lack interpretability, acting as black boxes that output only a score. Recently, MLLM-based methods(Zhou et al., 2025; Wang et al., 2025b; Mo et al., 2025; Gambashidze et al., 2025) have emerged, benefiting from the powerful multimodal understanding of MLLMs. For example, (Xu et al., 2024) and (Wang et al., 2025b) fine-tune MLLMs using large-scale visual question-answering pairs to achieve interpretable assessments. Nevertheless, most existing methods focus on general preferences and struggle with personalized preferences due to the scarcity of personalized data and complexity of individual tastes. Though (Salehi et al., 2024) stands out by proposing a personalized assessment method, it naively feeds reference and candidate images into an MLLM for score regression, thus lacking interpretability. In this work, we propose a reasoning-based personalized assessment system that achieves both accurate assessments and interpretable scoring.
@@
Image Preference Assessment Dataset is essential for training preference assessment models. (Xu et al., 2023) proposes ImageRewardDB, the first dataset for training models to assess general preferences (e.g. text-image alignment, aesthetics and safety), containing 137K candidate image pairs. Further, (Kirstain et al., 2023) and (Wu et al., 2023a) respectively developed PickaPic_v2 and HPD_v2, two large-scale and diverse datasets for learning human preferences. However, most existing public datasets focus on training models to capture general preferences, with limited consideration for personalized preferences. While (Salehi et al., 2024) curates a simulated personalized preference dataset, it remains unreleased. In this work, we construct a large-scale personalized assessment dataset annotated with diverse preference profiles and CoT-style reasoning to advance this field.
@@
Reasoning-based Multimodal Large Language Models have achieved significant progress in multimodal reasoning abilities through CoT-based or RL-based fune-tuning. Deepseek-R1 (Guo et al., 2025a) employs the GRPO algorithm (Shao et al., 2024) with the rule-based reward to improve reasoning abilities of LLM without critic models. Inspired by this, several works have incorporated GRPO-based post-training into MLLMs to enhance the multimodal reasoning in various tasks, including visual perception (Jiang et al., 2025; Huang et al., 2025; Liu et al., 2025b; Zhang et al., 2025a; Yu et al., 2025; Liu et al., 2025a; You & Wu, 2025) and visual understanding (Chen et al., 2025; Pan et al., 2025; Ni et al., 2025), math problem solving (Guo et al., 2025b; Zhang et al., 2025b; Li et al., 2025b), and visual quality assessment (Wu et al., 2025; Li et al., 2025a; Cai et al., 2025).
@@
Closest to ours, (Wang et al., 2025b) explores reasoning-based image preference assessment through RL-based post-training. However, this method focuses on assessing general preferences. Distinctly, PreferThinker is the first reasoning-based personalized assessment system. Inspired by Deepseek-R1, we adopt a two-stage training strategy comprising cold-start SFT and GRPO-based RFT to enable the system to master the reasoning capability Furthermore, we propose a similarity-aware prediction reward to better predict user’s preference profile, facilitating more reasonable evaluations exploration.
@@ 3.2 Large-scale CoT-style Personalized Preference Assessment Dataset
To enable the model to master CoT-style reasoning capabilities, we construct a large-scale CoT-style dataset that provides high-quality reasoning supervision. Since most existing public datasets focus on general preference assessment, we first construct a new dataset called PreferImg, dedicated to personalized image preference assessment, annotated with diverse visual preference profiles. We then employ the advanced Claude 3.7 model (Anthropic, 2024) to generate high-quality CoT-style annotations on the proposed PreferImg, ultimately producing our CoT-style dataset, PreferImg-CoT.
@@
Personalized Preference Assessment Data Generation. Due to the difficulty in acquiring personalized preference data with annotated preference profiles, such datasets are currently unavailable. To fill this gap, we construct a large-scale personalized assessment dataset containing 80K simulated users, each with a distinct preference profile. The data generation pipeline is shown in the Fig. 4(a). Specifically, we first randomly sample five visual preference elements to assign visual preference and non-preference profiles to each user. To account for the fact that real users may have multiple personalized preferences, several preference profiles are assigned to a subset of users. Then, we combine these profiles with initial prompts and feed into a text-to-image model to generate each user’s reference images (preferred and non-preferred) and two candidate images. To further guarantee content diversity, we select 190 K prompts from Lexica, DiffusionDB (Wang et al., 2022), and COCO (Lin et al., 2014) as initial prompts. Ultimately, we establish a dataset annotated with preference profiles, comprising 80 K users, 20K of whom are associated with multiple preference profiles, along with 1.36 million images, laying a solid foundation for subsequent CoT-style dataset construction.
@@ 3.3 Training Strategy
Stage 1: Supervised Fine-tuning for Cold-start Initialization. We utilize Qwen2.5-VL-7B (Bai et al., 2025) as the baseline model and conduct a SFT with the proposed CoT-style dataset PreferImg-CoT, as shown in Fig. 5(a). During the cold-start phase, the model learns to first predict the user’s preference profiles based on reference images and then leverages the predicted profiles as criteria to produce interpretable, multi-dimensional scores and assessments for candidate images. We train the model using a standard autoregressive language modeling objective and apply a token-level cross-entropy loss to provide strong supervision throughout the generation process:
@@
Stage 2: GRPO-based Reinforcement Learning for Post-training. Although cold-start endows the model with structured reasoning capabilities, its supervised training paradigm limits the exploration of diverse reasoning paths and restricts generalizability. To explore more reasonable assessments and enhance generalizability, we employ the Group Relative Policy Optimization (GRPO)-based reinforcement learning for exploration-driven post-training, as shown in Fig. 5(b). Unlike traditional policy optimization methods (Schulman et al., 2017), GRPO efficiently generates diverse reasoning responses by optimizing policy gradients on a sample group without requiring a critic model.
@@ 3.4 Similarity-aware Preference Profile Prediction Reward
where $Sim_{\text{text}}$ measures the similarity between two texts, $V_{+}$ and $V_{-}$ are ground-truth preference and non-preference profiles. We use SBERT (Reimers & Gurevych, 2019) to compute the text similarity.
Second, considering that image and text are the two sides of the same coin and should be tightly coupled. Therefore, we also measure the prediction accuracy via computing the image similarity $s_{\text{img}}$ between the generated images based on the predicted and ground-truth visual preference profiles. Specifically, given an initial text $T_{\text{initial}}$, we recap it by incorporating the predicted and ground-truth visual preference profiles, and then generate images using T2I model (Labs, 2024):
@@
where $Sim_{\text{img}}$ measures the similarity between two images. We use DreamSim (Fu et al., 2023) to compute the image similarity. Finally, the profile prediction reward function is a weighted combination of two similarity measures, designed to balance their contributions:
@@ 4.1 Experimental Settings
Datasets. We conduct experiments on the proposed PreferImg and real user dataset PickaPic (Kirstain et al., 2023). Based on the PreferImg, we construct a benchmark of 1,500 users, categorizing them into single-preference (SP) and multi-preference (MP) groups based on whether reference images reflect singular or multiple preferences. To evaluate generalization, we further divide the benchmark into seen and unseen subsets based on whether users’ profiles are included in the training data. Since no real-user dataset exists for personalized preference assessment, we process PickaPic, originally collected from real user interactions for general preference assessment, by grouping samples by user IDs, resulting in a user-specific benchmark with 894 user samples. Note that although the samples are grouped by users, the assessment labels still reflect general preferences, unlike PreferImg, which targets personalized preferences.
Implementation Details. PreferThinker is initialized from Qwen2.5-VL-7B and trained on a cluster of 8 NVIDIA A100 GPUs. We utilize the LLaMA-Factory (Zheng et al., 2024) for cold-start SFT, and train for only one epoch to prevent overfitting. After the SFT, we utilize the VLM-R1 framework (Shen et al., 2025) for RL-based post-training. We set a learning rate of 1e-5 with 6 rollout samples per input. The reward weights $w_{p}$, $w_{f}$ and $w_{a}$ are set to 0.7, 0.3 and 1.0. See Appendix A for details.
Baselines. We compare PreferThinker with (1) CLIP-based models: PickScore (Kirstain et al., 2023), ImageReward (Xu et al., 2023), HPSv2 (Wu et al., 2023a), CLIPScore (Radford et al., 2021), Aesthetics (Schuhmann et al., 2021), CycleReward (Bahng et al., 2025); (2) MLLM-based methods:, UnifiedReward (Wang et al., 2025c) UnifiedReward-Think (Wang et al., 2025b), LLaVA-Reward (Zhou et al., 2025), and ViPer (Salehi et al., 2024); (3) Open-source MLLMs: Qwen2.5-VL-7B (Bai et al., 2025), InternVL-3.5 (Wang et al., 2025a), GLM-4.1V-9B (Hong et al., 2025); (4) Closed-source MLLMs: Claude 3.7, GPT-4o, Doubao-1.5-vision, and Gemini-2.5. For a fair comparison, we categorize the methods based on whether support reference images as input for in-context learning.
@@ Appendix A Experimental Settings
We utilize Qwen2.5-VL-7B Instruct model as our base model and perform supervised fine-tuning using LLaMA-Factory (Zheng et al., 2024). We train the model on the proposed CoT-style cold-start dataset PreferImg-CoT for one epoch. The dataset includes 50,000 user samples with a single preference and 10,000 user samples with multiple preferences. we freeze the parameters of the vision tower and the multi-modal projector, and fine-tune the LLM. We employ DeepSpeed Zero-3 optimization strategy to handle the memory requirements of large models. The training settings of SFT are detailed in the Table S2.

### v2

@@ 1 Introduction
Image preference assessment is essential for both evaluating and aligning generative models with human preferences, playing a key role in applications like content recommendation. However, most existing works (kirstain2023pick; xu2023imagereward) focus on assessing general preferences, such as text-image alignment and aesthetics, relying on large-scale general preference data. The practical but challenging task of personalized preference assessment remains largely underexplored, where only a limited amount of personalized data is available for each user as prior information to assist in the assessment. As shown in Fig. 1 (a), personalized preference assessment is particularly challenging for two main reasons. First, unlike easily scalable general preference data, where most users share similar assessment criteria, each user’s personalized data is typically limited and distinct from others, making large-scale, user-specific training difficult. Second, personalized preferences are often complex and diverse, covering multiple dimensions such as art style, color, and art medium, which are more challenging than clear general human preferences (e.g. text-image alignment and aesthetics).
Current research on image preference assessment mainly falls into two categories: CLIP-based methods and Multimodal Large Language Model (MLLM)-based methods. CLIP-based methods (wu2023human; wu2023human2) typically fine-tune the CLIP model using large-scale text-image pairs with similarity to assess general preferences. However, they struggle with personalized preferences since the amount of personalized data is typically limited for large-scale training. Besides, simple similarity fails to capture complex individual preferences and lacks interpretability, as it outputs only a numerical score.
MLLM-based methods (wang2025unified; xu2024visionreward) achieve interpretable preference assessment by fine-tuning the MLLM on large-scale visual question-answering (VQA) pairs. However, the scarcity of personalized images makes it difficult to obtain sufficient VQA pairs for large-scale training, which limits their effectiveness for personalized preference assessment. Overall, most existing methods are designed for general preferences and therefore struggle to handle personalized preference. Recently, an MLLM-based method named ViPer (salehi2024viper) was proposed for personalized preference assessments. It directly feeds both personalized reference images and candidate images into an MLLM and fine-tunes the model using a supervised score regression strategy. Nevertheless, it only implicitly utilizes reference images to assist in evaluating the candidate images, failing to fully leverage the critical prior information they contain. Moreover, it lacks interpretable reasoning steps that explain how to produce its final score based on the individual reference images.
@@ 2 Related Work
Image Preference Assessment is crucial for both evaluating and aligning generative models with human preferences. While metrics such as FID (heusel2017gans) and CLIP scores (radford2021learning) assess image quality and text-image consistency, they are limited in capturing human preferences. To address this, most works (kirstain2023pick; liang2024rich; li2024stable; zhang2024learning; ba2025enhancing) employ large-scale preference data to fine-tune CLIP models for assessing general preferences. However, they lack interpretability, acting as black boxes that output only a score. Recently, MLLM-based methods(zhou2025multimodal; wang2025unifiedthink; mo2025learning; gambashidze2025listener) have emerged, benefiting from the powerful multimodal understanding of MLLMs. For example, (xu2024visionreward) and (wang2025unifiedthink) fine-tune MLLMs using large-scale visual question-answering pairs to achieve interpretable assessments. Nevertheless, most existing methods focus on general preferences and struggle with personalized preferences due to the scarcity of personalized data and complexity of individual tastes. Though (salehi2024viper) stands out by proposing a personalized assessment method, it naively feeds reference and candidate images into an MLLM for score regression, thus lacking interpretability. In this work, we propose a reasoning-based personalized assessment system that achieves both accurate assessments and interpretable scoring.
@@
Image Preference Assessment Dataset is essential for training preference assessment models. (xu2023imagereward) proposes ImageRewardDB, the first dataset for training models to assess general preferences (e.g. text-image alignment, aesthetics and safety), containing 137K candidate image pairs. Further, (kirstain2023pick) and (wu2023human2) respectively developed PickaPic_v2 and HPD_v2, two large-scale and diverse datasets for learning human preferences. However, most existing public datasets focus on training models to capture general preferences, with limited consideration for personalized preferences. While (salehi2024viper) curates a simulated personalized preference dataset, it remains unreleased. In this work, we construct a large-scale personalized assessment dataset annotated with diverse preference profiles and CoT-style reasoning to advance this field.
@@
Reasoning-based Multimodal Large Language Models have achieved significant progress in multimodal reasoning abilities through CoT-based or RL-based fune-tuning. Deepseek-R1 (guo2025deepseek) employs the GRPO algorithm (shao2024deepseekmath) with the rule-based reward to improve reasoning abilities of LLM without critic models. Inspired by this, several works have incorporated GRPO-based post-training into MLLMs to enhance the multimodal reasoning in various tasks, including visual perception (jiang2025rex; huang2025vision; liu2025visual; zhang2025improving; yu2025perception; liu2025seg; you2025seg) and visual understanding (chen2025chart; pan2025medvlm; ni2025point), math problem solving (guo2025decoupled; zhang2025r1; li2025vision), and visual quality assessment (wu2025visualquality; li2025q; cai2025q).
@@
Closest to ours, (wang2025unifiedthink) explores reasoning-based image preference assessment through RL-based post-training. However, this method focuses on assessing general preferences. Distinctly, PreferThinker is the first reasoning-based personalized assessment system. Inspired by Deepseek-R1, we adopt a two-stage training strategy comprising cold-start SFT and GRPO-based RFT to enable the system to master the reasoning capability Furthermore, we propose a similarity-aware prediction reward to better predict user’s preference profile, facilitating more reasonable evaluations exploration.
@@ 3.2 Large-scale CoT-style Personalized Preference Assessment Dataset
To enable the model to master CoT-style reasoning capabilities, we construct a large-scale CoT-style dataset that provides high-quality reasoning supervision. Since most existing public datasets focus on general preference assessment, we first construct a new dataset called PreferImg, dedicated to personalized image preference assessment, annotated with diverse visual preference profiles. We then employ the advanced Claude 3.7 model (anthropic2024claude) to generate high-quality CoT-style annotations on the proposed PreferImg, ultimately producing our CoT-style dataset, PreferImg-CoT.
@@
Personalized Preference Assessment Data Generation. Due to the difficulty in acquiring personalized preference data with annotated preference profiles, such datasets are currently unavailable. To fill this gap, inspierd by (salehi2024viper) and (anonymous2025prefgen), we construct a large-scale personalized assessment dataset containing 80K simulated users, each with a distinct preference profile. The generation pipeline is shown in the Fig. 4(a). Specifically, we first randomly sample five visual preference elements to assign visual preference and non-preference profiles to each user. To account for the fact that real users may have multiple personalized preferences, several preference profiles are assigned to a subset of users. Then, we combine these profiles with initial prompts and feed into a text-to-image model to generate each user’s reference images (preferred and non-preferred) and two candidate images. To further guarantee content diversity, we follow the PrefGen (anonymous2025prefgen) to select 190 K prompts from Lexica, DiffusionDB (wang2022diffusiondb), and COCO (lin2014microsoft) as initial prompts. Ultimately, we establish a dataset annotated with preference profiles, comprising 80 K users, 20K of whom are associated with multiple preference profiles, along with 1.36 million images, laying a solid foundation for subsequent CoT-style dataset construction.
@@ 3.3 Training Strategy
Stage 1: Supervised Fine-tuning for Cold-start Initialization. We utilize Qwen2.5-VL-7B (bai2025qwen2) as the baseline model and conduct a SFT with the proposed CoT-style dataset PreferImg-CoT, as shown in Fig. 5(a). During the cold-start phase, the model learns to first predict the user’s preference profiles based on reference images and then leverages the predicted profiles as criteria to produce interpretable, multi-dimensional scores and assessments for candidate images. We train the model using a standard autoregressive language modeling objective and apply a token-level cross-entropy loss to provide strong supervision throughout the generation process:
@@
Stage 2: GRPO-based Reinforcement Learning for Post-training. Although cold-start endows the model with structured reasoning capabilities, its supervised training paradigm limits the exploration of diverse reasoning paths and restricts generalizability. To explore more reasonable assessments and enhance generalizability, we employ the Group Relative Policy Optimization (GRPO)-based reinforcement learning for exploration-driven post-training, as shown in Fig. 5(b). Unlike traditional policy optimization methods (schulman2017proximal), GRPO efficiently generates diverse reasoning responses by optimizing policy gradients on a sample group without requiring a critic model.
@@ 3.4 Similarity-aware Preference Profile Prediction Reward
where $Sim_{\text{text}}$ measures the similarity between two texts, $V_{+}$ and $V_{-}$ are ground-truth preference and non-preference profiles. We use SBERT (reimers2019sentence) to compute the text similarity.
Second, considering that image and text are the two sides of the same coin and should be tightly coupled. Therefore, we also measure the prediction accuracy via computing the image similarity $s_{\text{img}}$ between the generated images based on the predicted and ground-truth visual preference profiles. Specifically, given an initial text $T_{\text{initial}}$, we recap it by incorporating the predicted and ground-truth visual preference profiles, and then generate images using T2I model (flux2024):
@@
where $Sim_{\text{img}}$ measures the similarity between two images. We use DreamSim (fu2023dreamsim) to compute the image similarity. Finally, the profile prediction reward function is a weighted combination of two similarity measures, designed to balance their contributions:
@@ 4.1 Experimental Settings
Datasets. We conduct experiments on the proposed PreferImg and real user dataset PickaPic (kirstain2023pick). Based on the PreferImg, we construct a benchmark of 1,500 users, categorizing them into single-preference (SP) and multi-preference (MP) groups based on whether reference images reflect singular or multiple preferences. To evaluate generalization, we further divide the benchmark into seen and unseen subsets based on whether users’ profiles are included in the training data. Since no real-user dataset exists for personalized preference assessment, we process PickaPic, originally collected from real user interactions for general preference assessment, by grouping samples by user IDs, resulting in a user-specific benchmark with 894 user samples. Note that although the samples are grouped by users, the assessment labels still reflect general preferences, unlike PreferImg, which targets personalized preferences.
Implementation Details. PreferThinker is initialized from Qwen2.5-VL-7B and trained on a cluster of 8 NVIDIA A100 GPUs. We utilize the LLaMA-Factory (zheng2024llamafactory) for cold-start SFT, and train for only one epoch to prevent overfitting. After the SFT, we utilize the VLM-R1 framework (shen2025vlm) for RL-based post-training. We set a learning rate of 1e-5 with 6 rollout samples per input. The reward weights $w_{p}$, $w_{f}$ and $w_{a}$ are set to 0.7, 0.3 and 1.0. See Appendix A for details.
Baselines. We compare PreferThinker with (1) CLIP-based models: PickScore (kirstain2023pick), ImageReward (xu2023imagereward), HPSv2 (wu2023human2), CLIPScore (radford2021learning), Aesthetics (schuhmann2021laion), CycleReward (bahng2025cycle); (2) MLLM-based methods:, UnifiedReward (wang2025unified) UnifiedReward-Think (wang2025unifiedthink), LLaVA-Reward (zhou2025multimodal), and ViPer (salehi2024viper); (3) Open-source MLLMs: Qwen2.5-VL-7B (bai2025qwen2), InternVL-3.5 (wang2025internvl3), GLM-4.1V-9B (hong2025glm); (4) Closed-source MLLMs: Claude 3.7, GPT-4o, Doubao-1.5-vision, and Gemini-2.5. For a fair comparison, we categorize the methods based on whether support reference images as input for in-context learning.
@@ Appendix A Experimental Settings
We utilize Qwen2.5-VL-7B Instruct model as our base model and perform supervised fine-tuning using LLaMA-Factory (zheng2024llamafactory). We train the model on the proposed CoT-style cold-start dataset PreferImg-CoT for one epoch. The dataset includes 50,000 user samples with a single preference and 10,000 user samples with multiple preferences. we freeze the parameters of the vision tower and the multi-modal projector, and fine-tune the LLM. We employ DeepSpeed Zero-3 optimization strategy to handle the memory requirements of large models. The training settings of SFT are detailed in the Table S2.